The coefficient of static friction between a sled and the snow is .15, with the coefficient of kinetic friction of .13. A 275N boy sits on the 200 N sled. How much force directed parallel to the horizontal surface is required to start the sled in motion? How much force is required to keep the sled in motion?
Answer #1

Total weight of boy and sled = 275 + 200 = 475 N
This is the weight which is equal to the Normal force on the combined system

Force required to start motion = µs N = µs * 475 = 0.15 * 475 N = 71.25 N

Force required to keep the sled in motion =μk N =  μk * 475 = 0.13 * 475 N = 61.75 N